40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6884     Entry time: Wed Jun 27 16:23:12 2012
Author: yuta 
Type: Update 
Category: IOO 
Subject: AS and REFL on AP table aligned 

I touched steering mirrors for AS and REFL at AP table.
AS beam and REFL beam now hits cameras at center and their respective PDs.

What I did:
  1. Aligned Y arm and X arm.

  2. Locked FPMI and aligned BS + X arm by minimizing ASDC (DC output of the AS55 PD, C1:LSC-ASDC_OUT reached ~ -1.43).

  3. Put -2V offset to the OMC stage 2 in yaw to avoid AS clipping. The offset is currently given by SRS DS345 on AUX_OMC_NORTH rack.

  4. Misaligned ETMs, locked MI in the bright fringe. Maximized ASDC (C1:LSC-ASDC_OUT reached ~ 1.22) by aligning 2 mirrors right after the vacuum chamber. This also centered beam spot on the AS camera.

  5. Locked MI in the dark fringe. Maximized REFLDC (DC output of the REFL55 PD, C1:LSC-REFLDC_OUT reached ~ 2.5) by aligning 2 mirrors after the vacuum chamber. Beam spot on the REFL camera was centered, too.

ELOG V3.1.3-